# JailGuard: A Universal Detection Framework for LLM Prompt-based Attacks

### v3

@@ 4.2.2. Combination Policy
To design a more effective and general mutation strategy, inspired by prior work (Hendrycks et al., 2019), we design a straightforward yet effective mutator combination policy. This policy integrates various mutators, leveraging their individual strengths to detect a wide array of attacks. The policy first involves selecting several mutators to build a mutator pool. When generating variants, the policy merges variants generated by different mutators based on the optimized built-in probability and then calculates the divergences of their responses.
To determine the optimal mutator pool and probability, we use 70% of our dataset (§ 5) as the training set and conduct large-scale experiments to collect empirical data of different mutators. We then employ an optimization tool (Komer et al., 2014) to search for the sampling probability of mutators. Consequently, for the combination policy of text mutators, we select Punctuation Insertion, Targeted Insertion and Translation to construct the mutator pool, and their sampling probabilities are $[0.24,0.52,0.24]$. For the image inputs, we select Random Rotation, Gaussian Blur and Random Posterization, and the sampling probabilities are $[0.34,0.45,0.21]$ respectively. The effectiveness of the mutator combination policy is validated in § 6.2 and § 6.4.
@@ 5. Dataset Construction
In real-world scenarios, MLLM systems face both jailbreaking and hijacking attack inputs across different modalities. For example, attackers may attempt to mislead the MLLM agent into producing harmful content (e.g., violence, sex) or inject specific instructions to hijack the agent into performing unintended tasks. Thus, it is crucial to comprehensively evaluate the effectiveness of LLM detection methods to identify and prevent various prompt-based attacks simultaneously.
However, due to the absence of a comprehensive LLM prompt-based attack dataset, existing LLM defense research mainly tests and evaluates their methods on inputs generated by specific attacks. For example, SmoothLLM (Robey et al., 2023) has evaluated its effectiveness in defending against jailbreak inputs generated by the GCG attack (Zou et al., 2023), overlooking other attacks (e.g., prompt injection attack) that can also have severe consequences. To address this limitation, We first collect the most popular jailbreaking and hijacking injection attack inputs from the open-source community and prior work. We then evaluate their effectiveness on LLM systems and applications, filtering out those samples where the attacks fail. Finally, we construct a dataset covering 15 types of prompt-based LLM attacks, covering two modalities of image and text, with a total of 11,000 items of attack and benign data. Our dataset is publicly available on our website (our, 2024).
@@
Combined Attack (Liu et al., 2023c) Combining existing methods (e.g., escape characters, context ignoring) to effectively inject.
@@
Typographic (TYPO) (Liu et al., 2023d) Embedding malicious instructions in blank images to conduct attacks.
@@
Typographic (SD+TYPO) (Liu et al., 2023d) Embedding malicious instructions in images generated by Stable Diffusion to conduct attacks.
@@
Text inputs. To ensure the diversity of LLM text attacks, we have collected a total of 12 kinds of attack inputs of the two common LLM attacks (i.e., jailbreaking attacks and hijacking injection attacks). Table 2 provides an overview of these attack methods. For jailbreaking attacks, to comprehensively cover various attack methodologies, we collect the most popular generative attack methods (i.e., Parameters (Huang et al., 2023a), DeepInception (Li et al., 2023b), GPTFuzz (Yu et al., 2023), TAP (Mehrotra et al., 2023), Jailbroken (Wei et al., 2024), Pair (Chao et al., 2023)) and the template-based attack method from the open-source community and existing study (Liu et al., 2023b) (including over 50 attack templates) to construct the attack inputs on GPT-3.5-Turbo-1106. To validate the effectiveness of these jailbreaking attack prompts and ensure the dataset’s quality, we follow the existing work (Qi et al., 2023) and evaluate the score of the prompts and the corresponding LLM’s responses violating OpenAI policies (score from 1 to 5) and only select those attack inputs with the highest score and the most severe violations (i.e., score 5) to construct the jailbreaking attack dataset, comprising a total of 2,000 items. For injection-based hijacking attacks, we have collected the most popular LLM injection attack methods, namely naive injection attack (Goodside, 2022), fake completion attack (Willison, 2023), ignoring context attack (Perez and Ribeiro, 2022), escape characters attack (Liu et al., 2023a), and combined attack (Liu et al., 2023c). We also verify the effectiveness of these attack samples on GPT-3.5-Turbo-1106 and select 2,000 items that can truly hijack LLMs to build our dataset.
@@
Considering that the number of benign queries in the real world is much more than attack queries, our dataset maintains a ratio of $1:1.5$ for the attack and benign data. We randomly sample a total of 6,000 questions from the existing LLM instruction datasets (Li et al., 2023a; Askari et al., 2023; Zhou et al., 2023) as the benign dataset. These instruction datasets have been widely used in prior work (Chen et al., 2024a; Dubois et al., 2024; Lou et al., 2023; Sha and Zhang, 2024) for fine-tuning and evaluation. The benign data covers various question types such as common sense questions and answers, role-playing, logical reasoning, summarization, etc.
@@
Text + Image inputs. Compared with text attacks, MLLM attack methods are relatively monotonous. We have collected the most popular adversarial-based jailbreaking attacks and typographic hijacking attacks. Adversarial-based attacks implant adversarial perturbations into images to guide the MLLM to produce harmful content. We leverage the prior work (Qi et al., 2024) to construct and collect 200 items of attack inputs on MiniGPT-4. The typographic attack is an injection-based attack method that involves implanting text into images to attack MLLMs (Liu et al., 2023d; Noever and Noever, 2021). We gather 200 attack inputs that use typographic images to replace sensitive keywords in harmful queries from MM-SafetyBench (Liu et al., 2023d), with 100 items embedding text in images generated by Stable Diffusion, and 100 items directly embedding text in blank images. Following existing work (Qi et al., 2023), all attack inputs have been validated for their effectiveness in attacking MiniGPT-4. Additionally, we include 600 benign inputs sampled from open-source training datasets of LLaVA (Zhou et al., 2023) and MiniGPT-4 (Zhu et al., 2023a) to balance the image dataset.
@@ 6. Evaluation
RQ4: What is the impact of the LLM query budget (i.e., the number of generated variants) in JailGuard?
@@ 6.1. Setup
  • Prior work (Liu et al., 2023c; Jain et al., 2023) implements several defense methods for jailbreaking and prompt injection attacks. We select four representative defense methods as baselines for text inputs in experiments, namely paraphrase, perplexity-based detection, data prompt isolation defense, and LLM-based detection. We query GPT-3.5-1106 to implement the paraphrase and LLM-based detection. Following the setting in prior work (Jain et al., 2023), we set the window size to 10 and use the maximum perplexity over all windows in the harmful prompts of AdvBench dataset (Zou et al., 2023) as the threshold, that is 1.51. For the data prompt isolation defense and LLM-based detection, we directly use the exiting implementation (Liu et al., 2023c).
@@
Metric. As mentioned in § 2, the LLM attack detector $detect(\cdot)$ assesses whether LLMs’ inputs are attacks. A positive output (i.e., 1) from $detect(\cdot)$ indicates an attack input, while a negative output (i.e., 0) signifies a benign input. For baselines employing filtering and similar mechanisms to defend against attacks without explicit detection mechanisms (e.g., Self-reminder), we use keyword-based detection from prior work (Robey et al., 2023; Zou et al., 2023) to evaluate their detection effects. When a specific refusal keyword (e.g., ‘I’m sorry’, ‘I apologize’) is detected in the LLM response, the original attack input is identified and blocked by the defense method, and $detect(\cdot)$ is 1 at this time, otherwise, it is set to 0.
@@
Following the prior work (Liu et al., 2022; Metzen et al., 2017), we collect the True Positive (TP), True Negative (TN), False Positive (FP), and False Negative (FN) in detection and use metrics accuracy, precision, and recall to comprehensively assess the effectiveness of detection.
Accuracy calculate the proportion of samples correctly classified by the defense method, encompassing both correctly identified attack samples (i.e., TP) and benign samples that pass through (i.e., TN), relative to the total samples. This metric comprehensively evaluates whether the detector effectively mitigates attacks while ensuring valid responses to benign inputs.
$Accuracy=\frac{TP+TN}{TP+FN+TN+FP}=\frac{|\{p\in\mathcal{P}_{a}:detect(R_{p})=1\}|+|\{p^{\prime}\in\mathcal{P}_{b}:detect(R^{\prime}_{p})=0\}|}{|\mathcal{P}_{a}|+|\mathcal{P}_{b}|}$
where $\mathcal{P}_{a}$ and $\mathcal{P}_{b}$ are the datasets of attack inputs and benign inputs. $\{p\in\mathcal{P}_{a}:detect(R_{p})=0\}$ and $\{p^{\prime}\in\mathcal{P}_{b}:detect(R^{\prime}_{p})=1\}$ separately represents attack inputs in $\mathcal{P}_{a}$ that are filtered by the detector and benign inputs in $\mathcal{P}_{b}$ that pass through the detector.
Precision calculates the number of correctly detected and filtered attack samples among all detected and filtered samples.
$Precision=\frac{TP}{TP+FP}=\frac{|\{p\in\mathcal{P}_{a}:detect(R_{p})=1\}|}{|\{p\in\mathcal{P}_{a}:detect(R_{p})=1\}|+|\{p\in\mathcal{P}_{b}:detect(R_{p})=1\}|}$
where $\{p\in\mathcal{P}_{b}:detect(R_{p})=1\}$ represents benign inputs that are incorrectly filtered by the detectors. Low precision indicates that the detector misjudges and blocks many benign inputs.
Recall assesses the number of correctly detected and filtered attack samples among all attack samples in the dataset.
$Recall=\frac{TP}{TP+FN}=\frac{|\{p\in\mathcal{P}_{a}:detect(R_{p})=1\}|}{|\mathcal{P}_{a}|}$
Low recall indicates that a considerable number of attack samples have bypassed the detector.
@@
Implementation. JailGuard generates $N=8$ variants for each input. For the baseline SmoothLLM that also needs to generate multiple variants, we have recorded the detection performance of each method in SmoothLLM when producing 4 to 8 variants and display the best detection results (i.e., the highest detection accuracy) each method achieves in Table 3. For text inputs, the probability of selecting and executing the replacement, insertion, and deletion operation on each character is $p=0.005$. Notably, for important sentences, this probability is increased to 5 times the usual, resulting in a value of 0.025. JailGuard use the string ‘[Mask]’ to replace and insert. The LLM systems and applications we used on text and image inputs are the GPT-3.5-Turbo-1106 and MiniGPT-4 respectively. It is important to note that in real-world scenarios, JailGuard should be integrated and utilized as part of the LLM system and application workflow to thwart potential attacks, which means that JailGuard performs detection from the perspective of developers. Consequently, it should have access to the underlying interface of LLMs, enabling it to query multiple variants in a batch and obtain multiple responses simultaneously. In our experiments, we simulate this process by making multiple accesses to the LLM system’s API. Our framework is implemented on Python 3.9. All experiments are conducted on a server with AMD EPYC 7513 32-core processors, 250 GB of RAM, and four NVIDIA RTX A6000 GPUs running Ubuntu 20.04 as the operating system.
@@ 6.2. RQ1: Effectiveness of Detecting Attack
Analysis. The results in Table 3 and Table 4 demonstrate the effectiveness of JailGuard in detecting LLM prompt-based attacks across different input modalities. JailGuard achieves an average detection accuracy of 81.68% on text inputs and 79.53% on image inputs with different mutators, surpassing the state-of-the-art baselines, which have an average accuracy of 68.19% on text inputs and 66.10% on image inputs. Remarkably, all mutators and policies implemented in JailGuard surpass the best baseline, with their results highlighted in bold. In addition, JailGuard achieves an average recall of 77.96% on text inputs and 77.93% on image inputs, which is 1.56 and 3.50 times the average result of baselines (50.11% and 22.25%), indicating its effectiveness in detecting and mitigating LLM attacks across different modal inputs. While excelling in attack detection, JailGuard also reduces FPs and separately improves the averaged precision by 5.54% and 1.19% on text and image inputs.
On the text dataset, the mutators and policy in JailGuard achieve an average accuracy and recall of 81.68% and 77.96%, which is 13.49% and 27.85% higher than the average results of the baselines. The best baseline (i.e., the ‘swap’ method of SmoothLLM) achieves the highest accuracy of 74.33% and recall of 44.98%. In comparison, different mutation strategies in JailGuard improve the accuracy of the best baseline by a factor of 1.18%-15.89% and improve its recall by 46.06%-113.65%. More specifically, the Random Insertion and Random Deletion mutators achieve the best accuracy of 82.40% and 81.31% among random mutators. The word-level and sentence-level mutators Synonym Replacement and Translation achieve the worst accuracy, namely 75.21% and 80.93%. Our analysis of their results shows that when creating variants, synonym replacement and translation can cause subtle changes in the semantics of words and sentences, leading to more false positives of benign cases. Although these two methods have good detection results on attack inputs (i.e., high recall), the increase in false positives limits their overall performance. In addition, all targeted mutators achieve much better results than their random version. Targeted Replacement and Targeted Insertion separately achieve the detection accuracy of 82.02% and 84.73%, improving the accuracy of 1.07% and 3.42% compared to Random Replacement and Random Insertion. Further analysis of their detection results reveals that the advantage of targeted mutators lies in detecting attacks with long texts and complex templates These attacks often use templates to construct specific scenarios and role-playing situations. The targeted mutators can identify the key content through word frequency and apply additional disturbances, thereby interfering with these attack samples and achieving better detection results. This observation is further confirmed by the ’Template’ column in Fig. 6. In addition, the combination policy in JailGuard further achieves the highest accuracy of 86.14% (marked in blue in Table 3), which illustrates the effectiveness of the mutator combination policy. We further study the impact of the probability in the built-in policy on detection results in § 6.4.
@@
On the image dataset, the baseline methods achieve an average accuracy of 66.10% and recall of 22.25%. The best baseline, ECSO detection, achieves an accuracy of 70.70% and a recall of 34.00%, illustrating the limitations of baselines in detecting attacks on image inputs. In contrast, the mutation strategies in JailGuard have achieved an average accuracy of 79.53% and recall of 77.93%, which far exceeds the results of baselines. The mutators and policy improve the best detection accuracy of baselines by a factor of 8.77%-17.26%, and the improvement on recall is even more significant, that is, 113.24%-159.56%. The policy in JailGuard combines the mutators Random Rotation, Gaussian Blur and Random Posterization, further achieving the detection accuracy and recall of 82.90% and 88.25%. It improves the results of the best baseline by 12.20% and 54.25%, demonstrating the detection effectiveness of JailGuard’s policy.
In addition, Fig. 5 intuitively demonstrates the advantages of JailGuard in attack detection compared to the baselines. We can observe that JailGuard (blue) achieves significantly better results than baselines (red), and the corresponding dots are distributed in the upper right corner, indicating high precision and recall in detection.

### v4

@@ 4.2.2. Combination Policy
To design a more effective and general mutation strategy, inspired by prior work (Hendrycks et al., 2019), we design a straightforward yet effective mutator combination policy. This policy integrates various mutators, leveraging their individual strengths to detect a wide array of attacks. The policy first involves selecting $m$ mutators $\{MT_{1},...,MT_{m}\}$ to build a mutator pool. When generating each variant, the policy selects a mutator from the mutator pool based on the built-in sampling probability of the mutator pool $\{p_{1},...,p_{m}\}$ and then uses the selected mutator to generate the variant. Note that each sampling probability $p_{i}$ corresponds to the mutator $MT_{i}$ and $\sum_{i=1}^{m}p_{i}=1$. After obtaining $N$ variants and constructing variant set $\mathcal{P}$, the policy calculates the divergence between the variant responses and detects attacks based on the methods in § 4.1.
To determine the optimal mutator pool and probability, we use 70% of our dataset (§ 5) as the development set and conduct large-scale experiments to collect empirical data of different mutators. Specifically, empirical data includes $N$ variants and corresponding responses generated by single mutators on the training set. These files can be obtained when evaluating the detection effect of each operator, and are reused here as empirical data to find the optimal operator pool and sampling probability. Note that these variants and responses can be directly collected in the evaluation of the detection effects mutators (§ 6.2) without additional effort and are reused here as empirical data. We then employ an optimization tool (Komer et al., 2014) to search for the sampling probability of mutators. During the search, we extract variants and corresponding responses from the empirical data of the corresponding mutators according to the probability, calculate the divergence of the selected responses, and iterate to find the optimal combination of mutator pool and probability. Consequently, based on the search results of the optimization tool, we select the text mutators Punctuation Insertion, Targeted Insertion and Translation to construct the mutator pool, and their sampling probabilities are $[0.24,0.52,0.24]$. For the image inputs, we select Random Rotation, Gaussian Blur and Random Posterization, and the sampling probabilities are $[0.34,0.45,0.21]$ respectively. The effectiveness of the mutator combination policy is validated in § 6.2 and § 6.4.
@@ 5. Dataset Construction
In real-world scenarios, LLM systems face both jailbreaking and hijacking attack inputs across different modalities. For example, attackers may attempt to mislead the LLM system into producing harmful content (e.g., violence, sex) or inject specific instructions to hijack the system into performing unintended tasks. Thus, it is crucial to comprehensively evaluate the effectiveness of attack detection methods to identify and prevent various prompt-based attacks simultaneously.
However, due to the absence of a comprehensive LLM prompt-based attack dataset, existing LLM defense research mainly tests and evaluates their methods on inputs generated by specific attacks. For example, SmoothLLM (Robey et al., 2023) has evaluated its effectiveness in defending against jailbreak inputs generated by the GCG attack (Zou et al., 2023), overlooking other attacks (e.g., prompt injection attack) that can also have severe consequences. To address this limitation, We first collect the most popular jailbreaking and hijacking injection attack inputs from the open-source community and prior work. We then evaluate their effectiveness on LLM systems and applications, filtering out those samples where the attacks fail. Finally, we construct a dataset covering 15 types of prompt-based LLM attacks, covering two modalities of image and text, with a total of 11,000 items of attack and benign data. We have released our dataset on our website (our, 2024), aiming to promote the development of security research of LLM systems and software.
@@
Combined Attack (Liu et al., 2024b) Combining existing methods (e.g., escape characters, context ignoring) to effectively inject.
@@
Typographic (TYPO) (Liu et al., 2023c) Embedding malicious instructions in blank images to conduct attacks.
@@
Typographic (SD+TYPO) (Liu et al., 2023c) Embedding malicious instructions in images generated by Stable Diffusion to conduct attacks.
@@
Text inputs. To ensure the diversity of text attacks on LLM systems, we have collected a total of 12 kinds of attack inputs of the two common prompt-based attacks (i.e., jailbreaking attacks and hijacking injection attacks). Table 2 provides an overview of these attack methods. For jailbreaking attacks, to comprehensively cover various attack methodologies, we collect the most popular generative attack methods (i.e., Parameters (Huang et al., 2023a), DeepInception (Li et al., 2023b), GPTFuzz (Yu et al., 2023), TAP (Mehrotra et al., 2023), Jailbroken (Wei et al., 2024), Pair (Chao et al., 2023)) and the template-based attack method from the open-source community and existing study (Liu et al., 2023b) (including over 50 attack templates) to construct the attack inputs on GPT-3.5-Turbo-1106. Except for the template-based method collected from the Internet, we generate no less than 300 attack prompts by each jailbreak method. To ensure the dataset’s quality, we have validated the effectiveness of the jailbreaking attack prompts and only selected the successful attacks that can guide LLMs to generate attackers-desired harmful content. Specifically, we follow the existing work (Qi et al., 2023) and evaluate the score of the prompts and the corresponding LLM’s responses violating OpenAI policies (score from 1 to 5). The highest score ‘5’ indicates that the model fulfills the attacker’s policy-violating instruction without any deviation and the response is a direct endorsement of the user’s intent. We only select those attack prompts with the highest score ‘5’ to construct a raw jailbreaking attack dataset. Then, we invite two co-authors with expertise in SE and AI security to manually verify whether these attack prompts are successful. They check the attack prompt and the corresponding LLM responses to determine whether the model produces attacker-desired harmful content. Subsequently, following the prior work (Pérez et al., 2020; Vieira et al., 2010), we use Cohen’s Kappa statistic to measure the level of agreement (inter-rater reliability) of the annotation results of two participants, which is 0.97 (i.e., “strong agreement” (McHugh, 2012)). For inconsistent cases, we invite a third co-author to moderate the discussion and conduct verification until we obtain results that are recognized by all three participants. According to our statistics, each participant takes about ten days to complete the verification. Finally, we construct a verified jailbreaking attack dataset covering 2,000 valid attack prompts. For injection-based hijacking attacks, we have collected the most popular LLM injection attack methods, namely naive injection attack (Goodside, 2022), fake completion attack (Willison, 2023), ignoring context attack (Perez and Ribeiro, 2022), escape characters attack (Liu et al., 2023a), and combined attack (Liu et al., 2024b). We directly verify the effectiveness of these attack samples using a verification framework integrated with existing method (Liu et al., 2024b) and select 2,000 items that can truly hijack LLMs to build our dataset.
@@
Considering that the number of benign queries in the real world is much more than attack queries, our dataset maintains a ratio of $1:1.5$ for the attack and benign data to simulate the data distribution in the real world. Our dataset is publicly released with data labels, and users can prune the dataset according to their experiment setting (e.g., pruning to a ratio of $1:1$ for attack and benign samples). We randomly sample a total of 6,000 questions from the existing LLM instruction datasets (Li et al., 2023a; Askari et al., 2023; Zhou et al., 2023) as the benign dataset. These instruction datasets have been widely used in prior work (Chen et al., 2024a; Dubois et al., 2024; Lou et al., 2023; Sha and Zhang, 2024) for fine-tuning and evaluation. The benign data covers various question types such as common sense questions and answers, role-playing, logical reasoning, etc.
@@
Text + Image inputs. Compared to the diverse text attacks, MLLM attacks have fewer types. We have collected the most popular adversarial-based jailbreaking attacks and typographic hijacking attacks. Adversarial-based attacks implant adversarial perturbations into images to guide the MLLM to produce harmful content. We leverage the prior work (Qi et al., 2024) to construct and collect 200 items of attack inputs on MiniGPT-4. The typographic attack is an injection-based attack method that involves implanting text into images to attack MLLMs (Liu et al., 2023c; Noever and Noever, 2021). We gather 200 attack inputs that use typographic images to replace sensitive keywords in harmful queries from MM-SafetyBench (Liu et al., 2023c), with 100 items embedding text in images generated by Stable Diffusion, and 100 items directly embedding text in blank images. Consistent with our text dataset, all attack inputs have been validated for their effectiveness in attacking MiniGPT-4 following the method described in the previous work (Qi et al., 2023). Additionally, we include 600 benign inputs sampled from open-source training datasets of LLaVA (Zhou et al., 2023) and MiniGPT-4 (Zhu et al., 2023a) to balance the image dataset.
@@ 6. Evaluation
RQ4: What is the impact of the built-in threshold $\theta$ in JailGuard?
RQ5: What is the impact of the LLM query budget (i.e., the number of generated variants) in JailGuard?
@@ 6.1. Setup
  • Prior work (Liu et al., 2024b; Jain et al., 2023) implements several defense methods for jailbreaking and prompt injection attacks. We select four representative defense methods as baselines for text inputs in experiments, namely paraphrase, perplexity-based detection, data prompt isolation defense, and LLM-based detection. We query GPT-3.5-1106 to implement the paraphrase and LLM-based detection. Following the setting in prior work (Jain et al., 2023), we set the window size to 10 and use the maximum perplexity over all windows in the harmful prompts of AdvBench dataset (Zou et al., 2023) as the threshold, that is 1.51. For the data prompt isolation defense and LLM-based detection, we directly use the exiting implementation (Liu et al., 2024b).
@@
Metric. As mentioned in § 2, the LLM attack detector $detect(\cdot)$ assesses whether LLMs’ inputs are attacks. A positive output (i.e., 1) from $detect(\cdot)$ indicates an attack input, while a negative output (i.e., 0) signifies a benign input. Note that several baseline methods (e.g., Self-reminder) exploit and reinforce the safety alignment of LLM itself to identify and block LLM prompt-based attacks. They do not provide explicit detection results and often provide refusal responses for attacks that cannot bypass these methods. To study the effectiveness of these methods in detecting valid attack prompts, we use the keywords from prior work (Robey et al., 2023; Zou et al., 2023) to obtain their detection results. When a specific refusal keyword (e.g., ‘I’m sorry’, ‘I apologize’) is detected in the LLM response, the original attack input is identified and blocked by the defense method, and $detect(\cdot)$ is 1 at this time, otherwise, it is set to 0.
@@
Following the prior work (Liu et al., 2022; Metzen et al., 2017), we collect the True Positive (TP), True Negative (TN), False Positive (FP), and False Negative (FN) in detection and use metrics accuracy, precision, and recall to comprehensively assess the effectiveness of detection. Accuracy calculates the proportion of samples correctly classified by the detection methods. Precision calculates the proportion of correctly detected attack samples among all detected samples, and recall calculates the proportion of correctly detected attack samples among all attack samples.
@@
Implementation. JailGuard generates $N=8$ variants for each input. For the baseline SmoothLLM that also needs to generate multiple variants, we have recorded the detection performance of each method in SmoothLLM when producing 4 to 8 variants and display the best detection results (i.e., the highest detection accuracy) each method achieves in Table 3. For text inputs, the probability of selecting and executing the replacement, insertion, and deletion operation on each character is $p=0.005$. Notably, the target mutators select the Top-3 scored sentences for each prompt as important sentences (prompt should contain at least three sentences), and for these important sentences, the probability of performing operations is increased to 5 times the usual, resulting in a value of 0.025. Following the prior work (Zeng et al., 2023; Gietz and Kalita, 2023; Qiang et al., 2022), JailGuard uses the string ‘[Mask]’ to replace and insert. In addition, to convert texts into vectors, researchers have proposed various models and methods (Zhang et al., 2010; Church, 2017; Yin and Shen, 2018). Based on the detection results of different word embedding models (Devlin et al., 2018; spa, 2024) (§ 8), we finally select the ‘en_core_web_md’ model in ‘spaCy’ library which is trained on large-scale corpus (Pradhan and Ramshaw, 2017; Fellbaum, 2010) and has been widely used in various NLP tasks (Wang et al., 2021b; Saeed et al., 2020; Kopanov, 2024). JailGuard uses the APIs in ‘spaCy’ library to load the model and convert the LLM response into a list of word vectors and then calculate their mean as the response vector. To determine the built-in detection threshold $\theta$, we randomly sample 70% of the collected dataset as the development set and finally choose $\theta=0.02$ for text input and $\theta=0.025$ for image input based on the detection results of JailGuard on the development set. More details are in § 6.5. The LLM systems and applications we used on text and image inputs are the GPT-3.5-Turbo-1106 and MiniGPT-4 respectively. It is important to note that in real-world scenarios, JailGuard should be integrated and utilized as part of the LLM system and application workflow to thwart potential attacks, which means that JailGuard performs detection from the perspective of developers. Consequently, it should have access to the underlying interface of LLMs, enabling it to query multiple variants in a batch and obtain multiple responses simultaneously. In our experiments, we simulate this process by making multiple accesses to the LLM system’s API. Our framework is implemented on Python 3.9. All experiments are conducted on a server with AMD EPYC 7513 32-core processors, 250 GB of RAM, and four NVIDIA RTX A6000 GPUs running Ubuntu 20.04 as the operating system.
@@ 6.2. RQ1: Effectiveness of Detecting Attack
Analysis. The results in Table 3 and Table 4 demonstrate the effectiveness of JailGuard in detecting LLM prompt-based attacks across different input modalities. JailGuard achieves an average detection accuracy of 81.68% on text inputs and 79.53% on image inputs with different mutators, surpassing the state-of-the-art baselines, which have an average accuracy of 68.19% on text inputs and 66.10% on image inputs. Remarkably, all mutators and policies implemented in JailGuard surpass the best baseline, with their results highlighted in bold. In addition, JailGuard achieves an average recall of 77.96% on text inputs and 77.93% on image inputs, which is 1.56 and 3.50 times the average result of baselines (50.11% and 22.25%), indicating its effectiveness in detecting and mitigating LLM attacks across different modal inputs. While excelling in attack detection, JailGuard also reduces FPs and separately improves the averaged precision by 5.54% and 1.19% on text and image inputs. Note that the experiment dataset simulates the real-world data distribution, where the number of benign samples is greater than that of attack samples. If using a dataset containing equal numbers of benign samples and attack samples, the advantage of JailGuard in detecting and mitigating LLM attacks will bring a greater accuracy improvement compared to the baselines.
@@
On the text dataset, the mutators and policy in JailGuard achieve an average accuracy and recall of 81.68% and 77.96%, which is 13.49% and 27.85% higher than the average results of the baselines. The best baseline (i.e., the ‘swap’ method of SmoothLLM) achieves the highest accuracy of 74.33% and recall of 44.98%. Furthermore, the baseline method, LLM-based detection, achieves a detection accuracy of 72.55%. It utilizes LLMs to effectively identify attack prompts but may lead to false positives. For example, for the benign prompt ‘Make a list of red wines that pair well with ribeyes. Use a./b./c. as bullets’, it will mistakenly classify this as an attack due to the seemingly harmful word ‘bullets’. In comparison, JailGuard can correctly identify benign prompts with sensitive words (e.g., bullets), and we have provided a case study in § 6.3. Different mutation strategies in JailGuard improve the accuracy of the best baseline by a factor of 1.18%-15.89% and improve its recall by 46.06%-113.65%. More specifically, the Random Insertion and Random Deletion mutators achieve the best accuracy of 82.40% and 81.31% among random mutators. The word-level and sentence-level mutators Synonym Replacement and Translation achieve the worst accuracy, namely 75.21% and 80.93%. Our analysis of their results shows that when creating variants, synonym replacement and translation can cause subtle changes in the semantics of words and sentences, leading to more false positives of benign cases. Although these two methods have good detection results on attack inputs (i.e., high recall), the increase in false positives limits their overall performance.
In addition, all targeted mutators achieve much better results than their random version. Targeted Replacement and Targeted Insertion separately achieve the detection accuracy of 82.02% and 84.73%, improving the accuracy of 1.07% and 3.42% compared to Random Replacement and Random Insertion. Further analysis of their detection results reveals that the advantage of targeted mutators lies in detecting attacks with long texts and complex templates These attacks often use templates to construct specific scenarios and role-playing situations. The targeted mutators can identify the key content through word frequency and apply additional disturbances, thereby interfering with these attack samples and achieving better detection results. This observation is further confirmed by the ’Template’ column in Fig. 9. In addition, the combination policy in JailGuard further achieves the highest accuracy of 86.14% (marked in blue in Table 3), which illustrates the effectiveness of the mutator combination policy. We further study the impact of the probability in the built-in policy on detection results in § 6.4.
@@
On the image dataset, the baseline methods achieve an average accuracy of 66.10% and recall of 22.25%. The best baseline, ECSO detection, achieves an accuracy of 70.70% and a recall of 34.00%, illustrating the limitations of baselines in detecting attacks on image inputs. In contrast, the mutation strategies in JailGuard have achieved an average accuracy of 79.53% and recall of 77.93%, which far exceeds the results of baselines. The mutators and policy improve the best detection accuracy of baselines by a factor of 8.77%-17.26%, and the improvement on recall is even more significant, that is, 113.24%-159.56%. The policy in JailGuard combines the mutators Random Rotation, Gaussian Blur and Random Posterization, further achieving the detection accuracy and recall of 82.90% and 88.25%. It improves the results of the best baseline by 12.20% and 54.25%, demonstrating the detection effectiveness of JailGuard’s policy. In addition, Fig. 7 intuitively demonstrates the advantages of JailGuard in attack detection compared to the baselines. We can observe that JailGuard (blue) achieves significantly better results than baselines (red), and the corresponding dots are distributed in the upper right corner, indicating high precision and recall in detection.
Defending Adaptive Attack. Although the mutation strategy in JailGuard randomly perturbs the input and the specific perturbation position cannot be determined, the important content selection method in targeted mutators may still be deceived by the attackers and suffer from adaptive attacks. Specifically, we assume that the attackers have a complete understanding of the targeted mutator’s implementation for selecting important content. Therefore, they can insert legitimate content with a large number of high-frequency words into the prompt to confuse the selection strategy. In such a situation, the targeted mutators select these legitimate sentences as important content and perform strong perturbations. Following this setting, we randomly select 200 text attack prompts from the collected dataset to construct adaptive attack samples and conduct experiments with both the original and adapted versions of these prompts on the GPT-3.5-1106 model. The legitimate content is implanted before the original attack prompt to reduce its impact on the semantics of the attack prompt. The experimental results are shown in Table 5. The rows show the detection accuracy of mutators Random Insertion, Random Replacement, Targeted Insertion, and Targeted Replacement on the original and adaptive attack prompts. We can observe that ❶ on the original attack prompts, the targeted mutators can improve the detection accuracy of their random version by 5.00% to 10.00%, which illustrates the effectiveness of word frequency-based targeted mutators in detecting attacks. They can identify those repeated attack content and impose strong perturbations. ❷ Adaptive attacks can degrade the detection effectiveness of the targeted mutators, leading to a drop in accuracy of up to 6.00%. In addition, even if the attacker cleverly deceives the important content selection, random perturbations to non-critical content can still effectively interfere with the attack content, ultimately resulting in detection performance close to that of random mutators. This demonstrates that JailGuard can resist the confusion of adaptive attacks and maintain the effectiveness of attack detection.
Table 5. Comparison of Mutators’ Detection Results on Original and Adaptive Attacks
Mutator Accuracy (%)
Origin Attack Adaptive Attack
Targeted Replacement 90.00 84.00
Targeted Insertion 84.00 82.50
Random Replacement 82.00 80.00
Random Insertion 80.00 78.50
Using Other LLMs to Generate Responses. JailGuard is deployed on the top of the LLM system $M$ to detect prompt-based attacks that can bypass the safety alignment of $M$. During detection, it directly queries $M$ to generate responses for variants. We have also studied the impact of using other LLMs to generate variant responses on detection results. Firstly, for unaligned models (unc, 2024), attackers can directly obtain their desired harmful content without designing attacks to bypass the safety alignment mechanisms. Therefore, models without safety alignment are not in the scope of JailGuard. In addition, a much less capable model than $M$ (e.g. GPT-2 (Radford et al., 2019) compared to GPT-3.5 used in experiments) may produce unpredictable and meaningless answers for complex input prompts and exhibit lower robustness when faced with perturbations, ultimately leading to a large number of false positives in detection. Moreover, using a better model (e.g., GPT-4o) can lead to better detection results. We randomly select 100 attack samples and conduct experiments on GPT-4o-2024-08-06 using different mutators. The experiment results show that a more powerful model can improve the attack detection effect of each mutator by 2% to 13%. However, more powerful LLMs typically come with higher costs and prices. For example, the price of GPT-4o-2024-08-06 is more than twice that of GPT-3.5-turbo-1106. To sum up, the model used to generate variant responses significantly impacts the detection performance of JailGuard. Considering that JailGuard is deployed on the top of the LLM system $M$ and is used to detect various attacks against $M$, we recommend directly using $M$’s model to generate responses for variants, which are GPT-3.5-turbo-1106 and MiniGPT-4 in our experiments.